# Where is the Super-Virial Gas? The Supply from hot inflows

### v1

@@ 3.4 Where is the super-virial gas coming from?
Now the question is what is the mechanism by which this gas is heated. As noted above, CR heating and feedback heating do not seem to be responsible for heating to super-virial temperatures in our simulations. Therefore, we track the host gas particles with initial temperature T K, which reached their maximum temperature at a simulation time of 1 Gyr. We calculate the past temperature, density, and spherical radius of all such gas particles (host and satellite). We find that there are three sets of gas particles that contribute to the super-virial phase, with two sets from the host and one from the satellites. One set has initial temperature similar to the virial gas K (set-1 hereafter) whereas other set has initial temperature of a cold gas K (set-2 hereafter). There are another set (set-3 hereafter), which are gas particles, initially coming from satellites and found to be at super-virial temperature. We plot median and dispersion of properties of all these particles for each cases. We list the properties of set-1, set-2 and set-3 particles below :
@@ 3.4.1 Set-1: Host gas particles, TK
In Figure 3, blue curves show the time evolution of the temperature (top row), density (2nd row), spherical radius (3rd row), and scale height (bottom row) of host gas particles with initial temperature T K, which reached their maximum temperature at a simulation time of 1 Gyr. These gas particles have initial temperature K in all the case except for high CR run where the initial temperature was K. This is because for high CR case, the initial halo has a lower thermal pressure, and in turn a lower temperature-halo. The Figure shows that gas heats up slowly over time and reaches over K at t=1 Gyr.
@@
In the second row of Figure 3, we plot past density of these super-virial gas particles. All of them start from , which is a typical density of the diffuse virial gas. Following their rise to super-virial temperatures at 1 Gyr, they cool down, reducing their density to . Again high CR run starts from the lower density of than the density of other runs, as we are tracking the gas particles from a larger radius as we mention in next paragraph. But eventually they reach the same final density as all the other runs. Therefore, there is almost a factor of 2000 and 1000 increase in density in high CR and all the other cases respectively.
In the third panel of Figure 3, we plot the time evolution of the position of these super-virial gas. All of them are falling from outer halo (50-80 kpc) into the central 10 kpc region. For all the other cases the particles fall from 40-50 kpc whereas for high CR run, the gas particles fall from little farther away, 80-90 kpc.
@@
As the gas particles fall inwards, they start heating up as this infall changes the gravitational potential energy, which can be converted to thermal energy and slowly heat up the gas. Until 0.9 Gyr, the temperature only increases to K. At 0.9 Gyr, these gas particles roughly reach 10 kpc which is the radius of Galactic disc. At that point, they get compressed by going from spherical to disc geometry, as shown by bottom-most row of Figure 3. And due to this compression, the gas is heated to over K and becomes super-virial gas. Therefore, it is clear that set-1 gas particles are infalling onto the disc and they are heated up by compression, and hence we see enhancement of density and temperature. Once they heat up to the super-virial phase, they cool down immediately because they become rotationally supported and radiate thermal energy away after they join the disc, as shown by the time evolution of the spherical radius , which does not decline and remains constant after Gyr. Please note that the smooth evolution of the density and temperature curves suggests the heating is compressive rather than shocks, as shown by tracking of individual particles in Figure 3.
@@ 3.4.2 Set-2: Host gas particle, TK
In the top row of Figure 4, we plot the time evolution of the temperature (top row), density (2nd row), and spherical radius (bottom row) of the host gas particles with initial temperature T K, which reached their maximum temperature at a simulation time of 1 Gyr. The gas particles in these sets are very few except for high CR case and absent in no feedback case. These gas particles get colder and hotter over time by fluctuating between K and K until 0.9 Gyr where the gas heats up and reaches over K at about 1 Gyr. These gas particles are probably heating from stellar feedback, hence are absent in no feedback case and are more in high CR case due to the additional heating from CR feedback by larger amount of CR.
In the middle row of Figure 4, we plot past density of these gas particles. All of them start from and finally become right before the heating up at 1 Gyr. Then their density again drops down to once they heated up by feedback. In bottom row of Figure 4, we plot the past position of these super-virial gas. All of them remain in the 10 kpc to 20 kpc region.
Thus, we see that these set-2 gas particles represent the cold ISM gas, which gets colder and denser over time and may form stars. Then they suddenly heat up by stellar feedback. However, set-2 gas particles are less than 0.01 times the set-1 gas particles. As we can see in Figure 5, the fraction of total supervirial gas host mass is due to feedback and due to compressive heating. This implies the mass in the supervirial phase as traced by set-2 gas particles is smaller by a factor of , hence set-2 gas particles contribute very little to super-virial gas mass budget.
@@ 3.4.3 Set-3: Particles Initially Coming From Satellites
There is another set of gas particles that can heat up to super-virial temperature, as shown in Figure 6. The gas from satellites gets stripped and falls inward. As they accrete onto the galactic disc, they heat up due to the gravitational potential energy loss and compression, just as the set-1 gas particles. We showed in the Figure 5, that the fraction in this set is initially zero but starts developing at the time-scale of satellite gas infall, around 1.2 Gyr. It starts with as low as 0.1% of the supervirial budget , but slowly goes up to even 10% at 1.5 Gyr. Therefore, at later time, the contribution from satellite gas phase can also contribute to supervirial gas phase significantly.
@@ 3.5 What distinguishes gas which heats up to super-virial temperatures?
Orange curves in Figure 3 plot the time evolution of the temperature (top row), density (2nd row), spherical radius (3rd row), and scale height (bottom row) of infalling virial gas particles that do not get heated up to super-virial temperatures. As we can see, they are also falling inwards, and as they do that, the temperature and density of the gas increases.
@@
In Figure 7, we investigate the distribution of probability distribution function of the angle () to the galaxy rotation axis at t = 0 Gyr (left panel) and at maximum temperature (right panel), for all infalling gas particles with a virial temperature beyond 20 kpc. In the Left panel, we can see the gas particles that heat up to super-virial temperatures are closer to the rotation axis, whereas the gas particles that remain at virial temperature are further from the rotation axis. It implies that infalling virial gas which originates from close to the rotation axis reaches higher temperatures. This is consistent with the behavior of rotating hot inflows (or ‘rotating cooling flows’) where the hot CGM phase inflows and spins up on a cooling timescale (Hafen et al., 2022; Stern et al., 2024). In these accretion flows gas has lower densities near the rotation axis, due to rotation support (see figure 4f in Stern et al. 2024). When the hot inflow approaches the disk, it transitions from a spherical geometry to a disk geometry. During this change in geometry, gas particles near the rotation axis suffer smaller radiative losses (which scale as density squared, see figure 4k there), and experience larger compressive heating rates (which scale inversely with density, see figure 4l there).
@@
Supporting the rotating hot inflow interpretation, the right panel of Figure 7 shows that the distribution when the temperature is maximal is close to 90 degrees, distinctly different from the initial distribution. This means that the super-virial gas particles have indeed transitioned from a spherical to a disk geometry. Further support for this interpretation is that the gas cools immediately after reaching super-virial temperatures (Figure 3), as expected in rotating hot inflows which cool immediately onto the ISM after transition into a disk geometry (Stern et al., 2024).
@@
Therfore, some of the virial temperature gas is infalling from outer CGM ( kpc) to the inner CGM ( kpc). All of this inflowing gas loses gravitational energy and is heated up by a factor of . When it reaches the inner CGM, the gas gets further heated up by another factor of due to compressive heating when the hot inflow changes from a spherical to a disc geometry. This change in geometry is instigated by rotation support in the hot inflow. Immediately thereafter, the hot inflow cools and joins the ISM disc. Specifically, inflowing virial gas closer to the rotation axis of the galaxy suffers less radiative cooling and more compressive heating due to lower gas density near the axis, and therefore heats up more.
3.6 What amount of infalling virial gas heats up to super-virial phase?
@@
To investigate how much virial gas is infalling, we calculate a cumulative distribution function (CDF) of the final position (at t= 1 Gyr) of the gas particles which starts at virial temperature beyond 20 kpc at t=0 Gyr. In the left panel of Figure 8, we show that only 10% gas particles enter 20 kpc at Gyr. With these gas particles which are 20 kpc at t=1 Gyr, we make a CDF of temperature distribution in right panel of Figure 8, where we see only 1-2% of these gas will heat up to super-virial temperatures.
@@ 3.7 Fate of this super-virial gas
This supervirial gas will also eventually cools down once it is supported by rotation, as we see in Figure 3, 4, and 6. In these figures, we denote the grey-shaded region to be the time evolution of the gas after they become super virial. We can see all the supervirial gas cools down in the later snapshot, consistent with the hot rotating inflow solution discussed in Stern et al. (2024). But the cooled gas will immediately be replenished by heating of continuously inflowing virial gas by the same process. Therefore, we see a constant amount of supervirial gas over time.
To summarize, we investigate mechanisms driving the heating of virial gas within a galaxy, finding that gas particles exhibit distinct heating behaviors based on their initial conditions and interactions. Set-1 gas particles, originating from the halo, closer to the rotation axis, heat up as they infall where the gas converts gravitational energy to heat mainly via compressive heating. This process causes gas infalling close to the rotation axis to reach super-virial temperatures just before cooling and joining the disk. In contrast, set-2 gas particles, likely representing the cold ISM gas, heat up sporadically, potentially due to stellar feedback, contribute minimally to the super-virial gas mass budget. Additionally, gas particles stripped from satellites and accreting onto the disc also heat up via compressive heating, further contributing to the super-virial phase. Some virial gas particles fail to heat up to supervirial temperature as they infall further from the rotation axis hence suffers more radiative cooling and less compressive heating.
@@ 4 Discussion, Limitations and Future work
We also investigate the temperature profile in a run where the virial temperature phase of the CGM is initialized with a cooling flow rather then hydrostatic equilibrium, using the cooling flow solution from Stern et al. (2019). We find that our cooling flow run also produces super-virial gas within 20 kpc and its mass is of the same order as the other runs (Figure 2). This makes sense since HSE initial conditions are expected to develop into a cooling flow within a cooling time (Stern et al., 2019).
@@
Bottom panel of Figure 3 and right panel of Figure 7 clearly indicates that the super-virial gas is extra-planar, just above the ISM disk. It will thus be very interesting to investigate future X-ray observations of edge-on galaxies, which we will discuss in the follow-up paper. In future work, we will also post process and produce observable like X-ray emission and absorption maps from the simulation snapshots. This will both facilitate direct comparison with observed maps, and assist in distinguishing between emission and absorption origins.
@@
It is fascinating to see that conventional CR/feedback heating do not produce this super-virial phase as anticipated, whereas gravitational and compressive heating is the dominant heating mechanisms in our idealized GIZMO simulations with FIRE stellar feedback model. However, this still depends on the feedback prescription used in the simulation. With a different feedback prescription (e.g., Smith et al., 2024b, a), one may get a different result. For example, the idealized simulation by the recent work of Vijayan & Li (2022) reproduced this temperature component, likely to be caused due to heating by stellar feedback, but with a very high star-formation rate unlike the MW. Another recent work by Singh Bisht et al. (2024) conclude that the super-virial gas is a result of the stellar feedback in their idealized simulation. Additionally, we do not have AGN feedback, that can have larger impact than stellar feedback in terms of heating the gas (e.g., Weinberger et al., 2023; Cochrane et al., 2023; Qutob et al., 2023; Su et al., 2024). For example, a recent work by Ramesh et al. (2023) also examines the physical properties of gas within the CGM surrounding 132 MW-like central galaxies at a redshift of z=0, using data from the cosmological simulation TNG50, which is part of the IllustrisTNG project. They found that energy from supermassive black hole (SMBH)-driven kinetic winds heats up the gas to super-virial temperatures (exceeding to K). Therefore, it would be useful to explore several different simulations and investigate the similar questions to compare. It will not only give us an overall picture of this newly-found super-virial phase, but also can constrain varying feedback mechanisms in different simulations.
Also, another limitations of our work is that it does not consider the galaxy in cosmological context. Our future works will explore the same questions with different cosmological simulations such as FIRE zoom-in runs (Hopkins et al., 2018b, 2023; Wellons et al., 2023), HESITA (Libeskind et al., 2020), TNG (Weinberger et al., 2017; Pillepich et al., 2018) and EAGLE (Crain et al., 2015; Schaye et al., 2015) simulations. These insights will deepen our understanding of galactic gas dynamics and provide valuable implications for broader inquiries into galaxy evolution and structure formation.

### v2

@@ 3.4 Where is the super-virial gas coming from?
Now the question is, what is the mechanism by which this gas is heated up? As noted above, CR heating and feedback heating do not seem to be responsible for heating to super-virial temperatures in our simulations. Therefore, we track the host gas particles with initial temperature T K, which reached their maximum temperature at a simulation time of 1 Gyr. We calculate the past temperature, density, and spherical radius of all such gas particles (host and satellite). We find that three sets of gas particles contribute to the super-virial phase, according to their initial properties, with two sets coming from the host and one from the satellites. The first set comes from the host with an initial temperature similar to T K (set-1 hereafter). The second set is also from the host with the initial temperature of a cold gas TK (set-2 hereafter). There is another set (set-3 hereafter), made of gas particles initially coming from satellites and found to be at super-virial temperature after Gyr. We track these particles and plot the median and dispersion of properties of all the trajectories of these particles for each case. We list the properties of set-1, set-2 and set-3 particles below.
@@ 3.4.1 Set-1: Host gas particles, TK
In Figure 3, blue curves show the time evolution of the temperature (top row), density (2nd row), and spherical radius (3rd row) of host gas particles with initial temperature T K, which reached their maximum temperature at a simulation time of 1 Gyr. The top panel shows that these gas particles roughly have an initial temperature K in all the cases except for the high CR run, where the initial temperature was K. This is because, for the high CR case, the initial halo has a lower thermal pressure and, in turn, a lower temperature halo. The gas heats up slowly over time and reaches over K at t=1 Gyr. In the third panel of Figure 3, we plot the time evolution of the position of the super-virial gas particles. All of them are falling from the outer halo (50-80 kpc) into the central 10 kpc region. For the high CR run, the gas particles fall from a little farther away, kpc, and for all the other cases, the particles fall from 40-50 kpc. The overall gas temperature of the CGM volume filling hot gas of the CGM in the high CR run is lower by a factor of two, as a result of our requirement to maintain HSE in the initial conditions. Therefore, the hot virial gas infalling from farther out than other runs can only heat up to the super-virial temperature, and hence, super-virial gas appears later in this run than other runs, as shown in the top-left panel of Figure 2.
@@
In the second row of Figure 3, we plot the past density of these super-virial gas particles. All of them start from , which is a typical density of the diffuse virial gas. As their temperature increases, their density also slowly increases over time as they fall inwards. Following their rise to super-virial temperatures at 1 Gyr, they cool down, increasing their density to . Again, the high CR run starts from a lower density of than the density of other runs, as we are tracking the gas particles from a larger radius, as we mention in the next paragraph. Eventually, they reach the same final density as all the other runs. Therefore, there is almost a factor of times increase in density while heated up, whereas times increase in density after they cooled down.
In the fourth panel of Figure 3, we show the trajectory of these particles in the plane. This figure shows that the particles start from higher and z and land at smaller and z, which is near the galactic disk.
@@
As the gas particles fall inwards, they start heating up as this infall changes the gravitational potential energy, which can be converted to thermal energy and slowly heat the gas. Until 0.9 Gyr, the temperature only increases to K. At 0.9 Gyr, these gas particles roughly reach 10 kpc, which is the radius of the galactic disc. At that point, they either get compressed by higher pressure and heated to over K or they can get shock-heated. Therefore, it is clear that set-1 gas particles are infalling onto the disc, and they are heated up by adiabatic compression and/or shocks, and hence we see enhancement of density and temperature. Once they heat up to the super-virial phase, they cool down immediately because they become rotationally supported and radiate thermal energy away after they join the disc, as shown by the time evolution of the spherical radius , which does not decline and remains constant after Gyr. Now the question is whether it is a result of a shock or of adiabatic compression. The initial smooth evolution of the density and temperature curves is indicative of the heating by compression rather than shocks, as shown by tracking of individual particles in Figure 3. To investigate this further, we track the radial velocity of these particles over time in the 5th panel of Figure 3. All the runs show fluctuation in velocity between super-sonic velocity km/sec and sub-sonic velocity. This implies that it is most probably a combination of shock (where super-sonic) and adiabatic compression (where sub-sonic). Additionally, the 6th row of Figure 3 shows the tracking of entropy for the hot inflowing particles that heat up to super-virial at 1 Gyr. It shows a roughly constant entropy throughout, with an entropy jump at 1 Gyr, where heating happens. Therefore, it implies that the earlier smooth temperature increase is due to the adiabatic compression of virial gas from infalling into a denser part of the galaxy. However, at a later time, when the entropy jump occurs, part of the heating originates from shock. Therefore, it is the combination of shock and adiabatic compression that heats the gas to super-virial temperatures. In conclusion, the infalling gas predominantly produces super-virial gas in our simulation. Hence, we also see more super-virial gas in the no feedback run as the feedback exits along the vertical axis of the galaxy and directly opposes the infalling gas. As mentioned before, our “no-feedback run” is performed mainly as a control experiment to see the effect of a pure cooling flow. It enables us to disentangle the net effect of the cooling flow from that of the stellar feedback in creating the super-virial gas. Also, the CR feedback can put additional resistance to the infalling gas, creating more fluctuations.
In the bottom-left panel of Figure 2, we show the time evolution of the net mass flux of the hot gas at 10 kpc. We find that the value of net mass influx of the hot gas is /yr, negative signs denoting inflow. However, the high-CR run shows a stronger inflow at a later time due to its lower CGM temperature. The inflow rate is roughly comparable to the average star formation rate (SFR) of the host galaxy, shown in the bottom-right panel of Figure 2. In simulations including stellar feedback, the system evolves toward a self-regulated quasi-steady state, where the SFR becomes comparable to the hot gas inflow rate. The typical SFRs lie in the range that is consistent with observed values in Milky Way-mass galaxies.
We also plot the subset of the trajectories of these particles in Figure A.1 to show that the median and dispersion capture the essence of the trajectories well. When we look at the individual particle trajectories, we see that most of the particles follow the above theory. In some cases, however, the particles are located near the disk, and they have cooled down. They are later lifted by feedback and heated to super-virial gas. Thus, the stellar feedback contributes to the super-virial phase for set-1 particles, but only marginally ( by mass).
@@ 3.4.2 Set-2: Host gas particle, TK
In the top row of Figure 4, we plot the time evolution of the temperature (top row), density (2nd row), and spherical radius (bottom row) of the host gas particles with initial temperature T K, which reached their maximum temperature at a simulation time of 1 Gyr. The gas particles in this set are very few, except for the high CR case, and are absent in the no feedback case. These gas particles get colder and hotter over time by fluctuating between K and K until 0.9 Gyr, where the gas heats up and reaches over K at about 1 Gyr. These gas particles are heated from stellar feedback and hence are absent in the no feedback case, and are more in the high CR case due to the additional heating from CR feedback by a larger amount of CR.
In the middle row of Figure 4, we plot the past density of these gas particles. All of them start from and finally become right before the heating up at 1 Gyr. Then their density again drops down to once they are heated up by feedback. In the bottom row of Figure 4, we plot the past position of this super-virial gas. All of them remain in the 10 kpc to 20 kpc region.
Thus, we see that these set-2 gas particles represent the cold ISM gas, which gets colder and denser over time and may form stars. Then they suddenly heat up by stellar feedback. However, set-2 gas particles are less than 0.01 times the set-1 gas particles. As we can see in Figure 5, the fraction of total super-virial gas host mass is due to feedback and due to compressive heating. This implies the mass in the super-virial phase, as traced by set-2 gas particles, is smaller by a factor of ; hence, set-2 gas particles contribute very little to the super-virial gas mass budget.
@@ 3.4.3 Set-3: Particles Initially Coming From Satellites
There is another set of gas particles that can heat up to super-virial temperature, as shown in Figure 6. The gas from satellites gets stripped and falls inward. As they accrete onto the galactic disc, they heat up due to the gravitational potential energy loss and compression, just as the set-1 gas particles. We showed in Figure 5 that the fraction in this set is initially zero but starts developing at the timescale of satellite gas infall, around 1.2 Gyr. It starts with as low as 0.1% of the super-virial budget but slowly goes up to even 10% at 1.5 Gyr. Therefore, at a later time, the contribution from the satellite gas phase can also contribute significantly to the super-virial gas phase.
@@ 3.5 What distinguishes gas which heats up to super-virial temperatures?
Orange curves in Figure 3 plot the time evolution of the temperature (top row), density (2nd row), and spherical radius (3rd row) of infalling virial gas particles that do not heat up to super-virial temperatures. As we can see, they are also falling inwards, and as they do that, the temperature and density of the gas increase slightly. We can see that the gas, which is heated to super-virial temperatures, settles to the disk at smaller radii than the gas that does not reach super-virial temperatures. Smaller radii imply higher densities, which suggests more compression. In the 4th panel of Figure A.1, we also see that the gas reaching the super-virial phase is closer to the rotation axis than the gas that fails to reach super-virial temperatures. We dig into this issue further and calculate the angle subtended by the particle position with the rotation axis () for these gas particles.
@@
In the top panel of Figure 7, we plot the probability distribution function of the angle () subtended by the particle position to the galaxy rotation axis at t = 0 Gyr (left panel) and at maximum temperature (middle panel), for all infalling gas particles with a virial temperature beyond 20 kpc. In the Left panel, we can see that the gas particles that heat up to super-virial temperatures are closer to the rotation axis, whereas the gas particles that remain at virial temperature are further from the rotation axis. It implies that infalling virial gas, which originates from close to the rotation axis, reaches higher temperatures. This is consistent with the behavior of rotating hot inflows (or ‘rotating cooling flows’) where the hot CGM phase inflows and spins up on a cooling timescale (Hafen et al., 2022; Stern et al., 2024). In these accretion flows, gas has lower densities near the rotation axis due to rotation support (see figure 4f in Stern et al. 2024). When the hot inflow approaches the disk, it transitions from a spherical geometry to a disk geometry. During this change in geometry, gas particles near the rotation axis suffer smaller radiative losses (which scale as density squared, see figure 4k there) and experience larger compressive heating rates (see figure 4l there).
@@
Supporting the rotating hot inflow interpretation, the middle panel of Figure 7 shows that the distribution, when the temperature is maximal, is close to 90 degrees, distinctly different from the initial distribution. This means that the super-virial gas particles have indeed transitioned from a spherical to a disk geometry. Further support for this interpretation is that the gas cools immediately after reaching super-virial temperatures (Figure 3), as expected in rotating hot inflows, which cool immediately onto the ISM after transition into a disk geometry (Stern et al., 2024).
In the right panel of Figure 7, we also present the probability distribution function of specific angular momentum j at t = 0 Gyr. We see that lower angular momentum gas that is near the rotation axis is forced to fall directly down the axis and heats up. On the other hand, gas further from the rotation axis with a larger j does not fall directly and is not heated up to high temperatures. Therefore, the origin of the gas heated to super-virial temperatures is largely influenced by our axisymmetric model, where the halo rotates around a central axis.
@@
In summary, some of the virial temperature gas is infalling from the outer CGM ( kpc) to the inner CGM ( kpc). All of this inflowing gas loses gravitational energy and is heated up by a factor of . When it reaches the inner CGM, the gas gets further heated up by another factor of due to the combination of compressive heating and shock heating when the hot inflow changes from a spherical to a disc geometry. This change in geometry is instigated by rotation support in the hot inflow. Immediately thereafter, the hot inflow cools and joins the ISM disc. Specifically, inflowing virial gas closer to the rotation axis of the galaxy suffers less radiative cooling and more compressive heating and shock heating due to lower gas density near the axis and, therefore, heats more.
3.6 What amount of infalling virial gas heats up to the super-virial phase?
@@
To investigate how much virial gas is infalling, we calculate a cumulative distribution function (CDF) of the final position (at t=1 Gyr) of the gas particles, which starts at virial temperature beyond 20 kpc at t=0 Gyr. In the left panel of the bottom row of Figure 7, we show that only 10% of gas particles enter 20 kpc at Gyr. With these gas particles, which are 20 kpc at t=1 Gyr, we make a CDF of temperature distribution in the middle panel of the bottom row of Figure 7, where we see only 2-3% of these gas will heat up to super-virial temperatures.
@@ 3.7 Fate of this super-virial gas
The super-virial gas will also eventually cool down once it is supported by rotation, as we see in Figure 3, 4, and 6. We can see all the super-virial gas cools down in the later snapshot after they reach the maximum temperature, consistent with the hot rotating inflow solution discussed in Stern et al. (2024). But the cooled gas will immediately be replenished by heating the continuously inflowing virial gas by the same process. Therefore, we see a constant amount of super-virial gas over time.
To summarize, we investigate mechanisms driving the heating of the virial gas within a galaxy, finding that gas particles exhibit distinct heating behaviors based on their initial conditions and interactions. Set-1 gas particles, originating from the halo, closer to the rotation axis, heat up as they infall, where the gas converts gravitational energy to heat mainly via a combination of compressive heating and shock heating. This process causes gas falling close to the rotation axis to reach super-virial temperatures just before cooling and joining the disk. In contrast, set-2 gas particles, likely representing the cold ISM gas, heat up sporadically, potentially due to stellar feedback, contributing minimally to the super-virial gas mass budget. Additionally, gas particles stripped from satellites and accreting onto the disc also heat up via compressive heating, further contributing to the super-virial phase. Some infalling virial gas particles fail to heat up to super-virial temperature as they fall further from the rotation axis, hence suffering more radiative cooling and less compressive heating.
@@ 4 Discussion, Limitations and Future work
We also investigate the temperature profile in a run where the virial temperature phase of the CGM is initialized with a cooling flow rather than hydrostatic equilibrium, using the cooling flow solution from Stern et al. (2019). We find that our cooling flow run also produces super-virial gas within 20 kpc, and its mass is of the same order as the other runs (Figure 2). This makes sense since HSE initial conditions are expected to develop into a cooling flow within a cooling time (Stern et al., 2019).
@@
The top panel of Figure 1 indicates that the super-virial gas is extra-planar, just above the ISM disk. It will thus be very interesting to investigate future X-ray observations of edge-on galaxies, which we will discuss in the follow-up paper. In future work, we will also post-process and produce observables like X-ray emission and absorption maps from the simulation snapshots. This will both facilitate direct comparison with observed maps and assist in distinguishing between emission and absorption origins.
@@
Effect of different feedback: It is fascinating to see that conventional CR/feedback heating does not produce this super-virial phase as anticipated, whereas gravitational and compressive heating is the dominant heating mechanism in our idealized GIZMO simulations with the FIRE stellar feedback model. However, this still depends on the feedback prescription used in the simulation. With a different feedback prescription (e.g., Smith et al., 2024b, a), one may get a different result. For example, the idealized simulation by the recent work of Vijayan & Li (2022) reproduced this temperature component, likely caused due to heating by stellar feedback, but with a very high star-formation rate, unlike the MW. Another recent work by Singh Bisht et al. (2024) concludes that the super-virial gas results from the stellar feedback in their idealized simulation. Additionally, we do not have AGN feedback, which can have a larger impact than stellar feedback in terms of heating the gas (e.g., Weinberger et al., 2023; Cochrane et al., 2023; Qutob et al., 2023; Su et al., 2024). For example, recent work by Ramesh et al. (2023) also examines the physical properties of gas within the CGM surrounding 132 MW-like central galaxies at a redshift of z=0, using data from the cosmological simulation TNG50, which is part of the IllustrisTNG project. They found that energy from the supermassive black hole (SMBH)-driven kinetic winds heats the gas to super-virial temperatures (exceeding to K). Therefore, it would be useful to explore several different simulations and investigate similar questions to compare. It will not only give us an overall picture of this newly found super-viral phase, but also constrain varying feedback mechanisms in different simulations.
Absence of cosmological context: Another limitation of our work is that it does not consider the galaxy in the cosmological context. The CGM of real galaxies is shaped by a complex cosmological history, involving non-steady accretion, feedback-driven baryon cycling, and chemical inhomogeneities. While our simulations do not include full cosmological evolution or cold-mode filamentary accretion, they are purposefully designed to isolate and study the behavior of hot-mode accretion in a controlled environment. Our initial conditions feature a simulation domain that extends to three times the virial radius, providing a natural reservoir of hot gas that can flow inward during the simulation. This setup captures gravitational infall of gas at roughly the virial temperature into the inner halo, our region of interest, over the timescale of the simulation. Importantly, our focus is not on cosmological accretion per se, but rather on how virial-temperature CGM gas infalls toward the galactic disk and gives rise to the super-virial phase observed in X-ray studies. We emphasize that we are not claiming this is the only source of the super-virial gas, but are testing it as one of the sources. We also argue that a steady cooling flow, although simplified, can describe the zeroth-order effect of the hot accreting gas in cosmological simulations (e.g., Sultan et al. 2024). While the absence of cold filaments and cosmological assembly history is a caveat, our goal is to disentangle the hot gas dynamics in isolation, which is difficult to achieve in fully cosmological simulations due to their inherent complexity.
In addition, our simulations adopt a simplified, axisymmetric, and isolated galaxy model to enable a controlled investigation of hot inflow and the resulting formation of super-virial gas. While this setup allows us to disentangle specific physical processes, it does come with limitations. In particular, our CGM is initialized as a smooth, hydrostatic, and baryon-rich medium. Despite uncertainties remaining in the baryonic fraction of Milky Way-mass galaxies, it may have a higher baryon fraction than the Milky Way (Miller & Bregman, 2013, 2015) due to the absence of prior baryon loss from feedback or cosmological assembly. As this CGM cools, it naturally begins to collapse and drive inflows, with the infall rate evolving non-monotonically over time. This idealized setup may enhance infall and subsequent heating, potentially exaggerating the generation of super-virial gas in the early phases. As for the axial symmetry, while the net angular momentum assumed in our initial conditions is broadly consistent with full cosmological simulations (DeFelippis et al., 2020) and should capture the zeroth-order inflowing angular momentum of hot accreting gas, it may also result in a more coherent inflow of gas than occurs in reality. While we find qualitative similarities with expectations from cosmological simulations, such as feedback-driven regulation and transient heating, we acknowledge that the detailed evolution is sensitive to the initial conditions and timing of measurement. A fully cosmological context will be necessary to test the robustness of these results, which we leave for future work. Our future works will explore the same questions with different cosmological simulations, such as FIRE zoom-in runs (Hopkins et al., 2018b, 2023; Wellons et al., 2023), HESITA (Libeskind et al., 2020), TNG (Weinberger et al., 2017; Pillepich et al., 2018), and EAGLE (Crain et al., 2015; Schaye et al., 2015) simulations. These insights will deepen our understanding of galactic gas dynamics and provide valuable implications for broader inquiries into galaxy evolution and structure formation.
Effect of ad-hoc satellite distribution: The satellite configuration was originally designed to see satellite contribution to the super-virial gas, and it was found to be insignificant. However, such exploration is still important since we covered a wide parameter space of satellite mass and orbit, and still did not see a huge effect. The timescale of Gyr, when satellites start to contribute to super-virial gas, is indeed a product of the initial conditions. The infall time scale of these satellites from 100-150 kpc is roughly Gyr. A different satellite distribution in initial conditions can potentially make their effect important according to their infall time.
In summary, while our idealized simulations lack a full cosmological context, they provide a valuable framework for isolating and understanding the physical processes that give rise to super-virial gas in a controlled environment. This approach allows us to disentangle the roles of hot inflow, feedback, and gravitational dynamics without the added complexity of ongoing structure formation. Additionally, our exploration of satellite contributions sets the stage for future work using cosmological simulations, where such interactions will occur more naturally. Finally, this setup offers a useful testbed for examining the efficacy of FIRE stellar feedback in driving thermal evolution of the CGM; notably, we find that the feedback alone does not appear sufficient to generate significant super-virial heating in the absence of additional infall-driven processes.